(−2, −5), x − 2y = 3
Write an equation of the line containing the specified point and perpendicular to the indicated line



Answer by mananth(16946)

1 x -2 y = 3
Find the slope of this line

-2 y = -1 x + 3
Divide by -2
y = 0.50 x + -1.5
Compare this equation with y=mx+b
slope m = 1/2

The slope of a line perpendicular to the above line will be the negative reciprocal -2
m1*m2=-1
The slope of the required line will be -2

m= -2 ,point ( -2 , -5 )
Find b by plugging the values of m & the point in
y=mx+b
-5 = 4.00 + b
b= -9
m= -2
The required equation is y =-2x-9
m.ananth@hotmail.ca